# Preconditioned Score-based Generative Models

### v2

@@ Abstract
Score-based generative models (SGMs) have recently emerged as a promising class of generative models. However, a fundamental limitation is that their sampling process is slow due to a need for many (e.g., $2000$) iterations of sequential computations. An intuitive acceleration method is to reduce the sampling iterations which however causes severe performance degradation. We assault this problem to the ill-conditioned issues of the Langevin dynamics and reverse diffusion in the sampling process. Under this insight, we propose a model-agnostic preconditioned diffusion sampling (PDS) method that leverages matrix preconditioning to alleviate the aforementioned problem. PDS alters the sampling process of a vanilla SGM at marginal extra computation cost, and without model retraining. Theoretically, we prove that PDS preserves the output distribution of the SGM, no risk of inducing systematical bias to the original sampling process. We further theoretically reveal a relation between the parameter of PDS and the sampling iterations, easing the parameter estimation under varying sampling iterations. Extensive experiments on various image datasets with a variety of resolutions and diversity validate that our PDS consistently accelerates off-the-shelf SGMs whilst maintaining the synthesis quality. In particular, PDS can accelerate by up to 29×29$\times$29 × on more challenging high resolution (1024$\times$1024) image generation. Compared with the latest generative models (e.g., CLD-SGM, DDIM, and Analytic-DDIM), PDS can achieve the best sampling quality on CIFAR-10 at a FID score of 1.99. Our code is made publicly available to foster any further research https://github.com/fudan-zvg/PDS.
@@ 1 Introduction
As an alternative approach to generative adversarial networks (GANs) [1], recent score-based generative models (SGMs) [2, 3, 4, 5] have demonstrated excellent abilities in data synthesis (especially in high resolution images) with easier optimization [2], richer diversity [6], and more solid theoretic foundation [7]. Based on these progresses, SGMs and related models show great potential in various application[8] including high fidelity audio streams generation [9, 10], text-to-image generation [11, 12, 13], photo-realistic images editing [14, 15, 16, 17], video [18] generation, 3D shape generation [19], and scene graphs generation [8]. Starting from a sample initialized with a Gaussian distribution, a SGM produces a target sample by simulating a sampling process, typically a reverse diffusion, Langevin dynamics, or their combination (Alg. 1). Compared to the state-of-the-art GANs [20, 21, 22], a significant drawback with existing SGMs is drastically slower generation due to the need of taking many iterations for a sequential sampling process [4, 11, 6].
@@
Conceptually, the Langevin dynamics as defined in Eq. (1) can be considered as a special case of Metropolis adjusted Langevin algorithm (MALA) [23, 24, 25]. When the coordinates of the target dataset (e.g., the pixel locations of a natural image) have a widely various scales or are strongly correlated, the sampling process will be ill-conditioned. As a result, the convergence of the sampling process is not guaranteed when the step sizes are large [25]. This explains the failure of sampling under a direct acceleration (Fig. 1). Observing that the reverse diffusion has a similar update rule as the Langevin dynamics (Eq. (1) vs. Eq. (3)), we conceive that it suffers the same ill-conditioned issue.
——————$T=2000$——————$T=200$——————$T=133$——————$T=100$——————$T=66$
———–      Ours                 Baseline [4]
Figure 1: Facial images at a resolution of $1024\times 1024$ generated by NCSN++ [4] under a variety of sampling iterations (top) without and (bottom) with our PDS. It is evident that NCSN++ decades quickly with increasingly reduced sampling iterations, which can be well solved with PDS. In terms of running speed for generating a batch of 8 images, PDS reduces the time cost of generating $8$ human facial samples from $1920$ sec (more than half an hours) to $68$ sec (about one minute) on one NVIDIA RTX 3090 GPU. Dataset: FFHQ [21]. More samples in Appendix.
@@
In light of this insight as above, we propose an efficient, model-agnostic preconditioned diffusion sampling (PDS) method for accelerating existing pretrained SGMs without the need for model retraining. The key idea is to make the scales of the target distribution more similar along all the coordinates [23, 26] using a matrix preconditioning operation, hence solving the ill-conditioned problem of both the Langevin dynamics Eq. (1) and the reverse diffusion Eq. (3), as demonstrated in Fig. 2. Formally, we enrich the Langevin dynamics Eq. (1) by reformulating the sampling process as
@@
Theoretically, we prove that PDS can preserve both the steady-state distribution of the original Langevin dynamics and the final-state distribution of the original reverse diffusion, with the assistance of Fokker-Planck equation [27] and related techniques. We construct the preconditioning operator $M$ using Fast Fourier Transform (FFT) [28], making its computational cost marginal.
In this work, we make the following contributions: (1) For sampling acceleration, we introduce a novel preconditioned diffusion sampling (PDS) process. PDS preconditions the existing sampling process for alleviating ill-conditioned issues, with theoretical guarantee of keeping the original target distributions in convergence. (2) With PDS, a variety of pretrained SGMs can be accelerated significantly for image synthesis at various spatial resolutions, without model retraining. In particular, PDS delivers 29×29$\times$29 × reduction in wall-clock time for high-resolution $(1024\times 1024)$ image synthesis. Compared with latest alternatives (e.g., CLD-SGM [29], DDIM [30], and Analytic-DDIM [31]), PDS achieves superior performance on CIFAR-10 (FID score 1.99).
A preliminary version of this work was presented in ECCV 2022 [32]. In this paper, we have further extended our conference version as follows: (i) We reveal that the sampling process of the original SGMs is insensitive to the structural data (Sec. 3.4); (ii) We provide more theoretical explanation including its effect of preserving the final-state distribution of the reverse diffusion during sampling (Sec. 4.3); (iii) We conduct more quantitative experiments, including an extra dataset (CelebA), more iteration budgets evaluation, and comparing with more acceleration methods in both SGMs and DDPMs; (iv) We investigate the parameter space of PDS more comprehensively, resulting in further performance improvement; (v) We develop a theory easing the process of estimating proper parameter (Sec. 6.3) with experimental validation (Sec. 7.3).
@@ 2 Related work
Sohl-Dickstein et al. [33] first proposed to sequentially corrupt a target data distribution with Gaussian noises and train a model to learn the backward process to recover the target data distribution, inspired by non-equilibrium statistical physics. Exploring this procedure later on, Song and Ermon [2] proposed first SGMs, the noise conditional score network (NCSN). They [3] further proposed NCSNv2 for higher resolution image generation (e.g., $256\times 256$) by scaling noises and improving stability with moving average. Summarizing all the previous SGMs into a unified framework based on the stochastic differential equation (SDE), Song et al. [4] proposed NCSN++ to generate high-resolution images via numerical SDE solvers for the first time. Bortoli et al. [7] provided the first quantitative convergence results for SGMs. Vahdat et al. [34] trained SGMs in a latent space with the variational autoencoder framework. As another class of relevant generative models, denoising diffusion probabilistic models (DDPMs) [35, 36, 37, 11] also demonstrate excellent performance on image synthesis. They are mostly trained by maximizing an evidence lower bound (ELBO).
@@
There are some recent works on improving the sampling quality of SGMs and DDPMs. Song et al. [30] proposed DDIMs by improving the sampling of DDPMs via a non-Markovian sampling process. Liu et al. [38] further improved DDIMs by implementing the sampling process near a manifold to guarantee that the model can fit the ground-truth score function well. Kingma et al. [39] proposed variations diffusion models to learn the noise schedule during training with another neural network for higher log-likelihood of the sampling results. Using the Hamiltonian Monte Carlo methods [40], Dockhorn et al. [29] extended the sample space of SGMs with a velocity vector, and proposed critically-damped Langevin diffusion (CLD) based SGMs. Jolicoeur-Martineau et al. [41] used a numerical SDE solver with adaptive step sizes to accelerate SGMs. A number of recent works [42, 43, 44] have utilized the concept of knowledge distillation to expedite the sampling process of DDPMs. Their common idea is a progressive distillation approach, involving the iterative refinement of a pretrained model. In comparison, our proposed PDS comes with a unique advantage of not requiring to train additional models for sampling acceleration.
@@
As our PDS, some prior methods utilize non-isotropic noises to accelerate the sampling process. Jiang et al.[45] restricted the sampling process in a linear subspace where the data reside to accelerate the convergence of the sampling process. This might help with the illcondition issue and convergence speed, as the space of sampling decreases. However, our PDS directly adjusts the scales of different coordinates of the sampling space for alleviating ill-condition issue, without the need for subspace restriction. Nichol and Dhariwal [46] learned an additional variance vector during training a DDPM. Instead we exploit a preconditioning matrix simply estimated from the training data, without the need of training extra parameters. With diagonal covariance matrices, Bao et al. [31, 47] given the analytical formulation of a sequential of diagonal covariance matrices that maximizes the ELBO between the corrupting process and the sampling process for better sample performance. In comparison, our method is more efficient without solving a complex optimization problem.
@@
However, these above methods are limited in the following aspects: (1) They tend to introduce much extra computation. For example, CLD-SGMs double the dimension of data due to learning the velocity of the diffusion. Jolicoeur-Martineau et al. [41] added a high-order numerical solver that increases the number of calling the SGM, resulting in much more cost. With our PDS, the only extra calculation relates the preconditioning matrix that can be efficiently implemented by Fast Fourier Transform. As a result, PDS yields marginal extra computation to the vanilla SGMs. (2) They often improve a single specific SGM/DDPM while our PDS is able to accelerate three different SGMs (NCSN [2], NCSNv2 [3], and NCSN++ [4]). (3) Unlike our work, none of previous methods has demonstrated the scalability to more challenging high-resolution image generation tasks (e.g., FFHQ facial images with $1024\times 1024$ resolution).
@@ 3.1 Score matching
Assume that we observe i.i.d. samples from an unknown distribution $p_{\mathbf{x}^{\ast}}(\mathbf{x})$ of random vector $\mathbf{x}\in\mathbb{R}^{d}$, parameterized as $p_{\mathbf{x}^{\ast}}(\cdot;\theta)$ with $\theta$ the learnable parameters. In a non-normalized statistical learning setting, we have an explicit expression for $p_{\mathbf{x}^{\ast}}(\cdot;\theta)$ up to a normalization factor $Z(\theta)$
$p_{\mathbf{x}^{\ast}}(\mathbf{x};\theta)=\frac{1}{Z(\theta)}q(\mathbf{x};\theta),\forall\mathbf{x}\in\mathbb{R}^{d}$ (6)
@@
Score matching is an effective approach to solving this challenge [48]. Specifically, a model $\mathbf{s}_{\theta}(\cdot)$ is trained to directly approximate the score function $\mathbf{s}_{\mathbf{x}^{\ast}}(\cdot):=\bigtriangledown_{\mathbf{x}}\log p_{\mathbf{x}^{\ast}}(\cdot)$ through the objective
@@
At the absence of $Z(\theta)$, training with this objective is more tractable. [48] further proves that under some mild conditions, if $p_{\mathbf{x}^{\ast}}(\cdot)=p_{\mathbf{x}^{\ast}}(\cdot;\theta^{\ast})$, the ground-truth parameters $\theta^{\ast}$ can be achieved by minimizing Eq. (7) to zero.
However, Eq. (7) can not be evaluated in more general cases where an expression of the score function about $\theta$ is absent. To address this, [49] proposes an equivalent, simpler objective inspired by denoising autoencoders
@@
where $p_{\sigma}(\mathbf{x}_{\sigma}|\mathbf{x})=\mathcal{N}(\mathbf{x}_{\sigma};\mathbf{x},\sigma^{2}I)$. Formally, given a noised data sample $\mathbf{x}_{\sigma}$, the model $\mathbf{s}_{\theta}(\cdot)$ is trained to extract the Gaussian noise vector added into the sample. This denoises the sample corrupted by the noise, i.e., subtracting the noise vector gives the original sample. It is proved that minimizing the objective Eq. (8) allows the model to fit the ground-truth score function $\mathbf{s}_{\mathbf{x}^{\ast}}$ of the data distribution [49].
3.2 Sampling from the Langevin dynamics
SGMs generate samples from a target distribution through iteratively evaluating the target score function. The first SGM, noise conditional score network (NCSN) [2] $\mathbf{s}_{\theta}$, is trained on a series of score matching tasks via
$\sum_{t=1}^{T}\sigma^{2}_{t}\mathbb{E}_{\mathbf{x}\sim p_{\mathbf{x}}}\left[\left\|\mathbf{s}_{\theta}(\mathbf{x}_{\sigma_{t}},\sigma_{t})-\bigtriangledown_{\mathbf{x}_{\sigma_{t}}}\log p_{\sigma_{t}}(\mathbf{x}_{\sigma_{t}}|\mathbf{x})\right\|^{2}\right],$ (9)
with the standard deviations $\{\sigma_{t}\}_{t=1}^{T}$ satisfies $0<\sigma_{1}<\sigma_{2}<\ldots<\sigma_{T}$. At inference, a NCSN generates samples from $p_{\mathbf{x}^{\ast}}$ through the discretized Langevin dynamics

### v3

@@ Abstract
Score-based generative models (SGMs) have recently emerged as a promising class of generative models. However, a fundamental limitation is that their sampling process is slow due to a need for many (e.g., $2000$) iterations of sequential computations. An intuitive acceleration method is to reduce the sampling iterations which however causes severe performance degradation. We assault this problem to the ill-conditioned issues of the Langevin dynamics and reverse diffusion in the sampling process. Under this insight, we propose a novel preconditioned diffusion sampling (PDS) method that leverages matrix preconditioning to alleviate the aforementioned problem. PDS alters the sampling process of a vanilla SGM at marginal extra computation cost and without model retraining. Theoretically, we prove that PDS preserves the output distribution of the SGM, with no risk of inducing systematical bias to the original sampling process. We further theoretically reveal a relation between the parameter of PDS and the sampling iterations, easing the parameter estimation under varying sampling iterations. Extensive experiments on various image datasets with a variety of resolutions and diversity validate that our PDS consistently accelerates off-the-shelf SGMs whilst maintaining the synthesis quality. In particular, PDS can accelerate by up to 28×28$\times$28 × on more challenging high-resolution (1024$\times$1024) image generation. Compared with the latest generative models (e.g., CLD-SGM and Analytic-DDIM), PDS can achieve the best sampling quality on CIFAR-10 at an FID score of 1.99. Our code is publicly available to foster any further research https://github.com/fudan-zvg/PDS.
Keywords:
Image synthesis, Score-based generative models, Preconditioning
journal: International Journal of Computer Vision
@@ 1 Introduction
As an alternative approach to generative adversarial networks (GANs) (Goodfellow et al., 2014), recent score-based generative models (SGMs) (Yang et al., 2023) have demonstrated excellent abilities in data synthesis (especially in high-resolution images) with easier optimization (Song and Ermon, 2019), richer diversity (Xiao et al., 2022), and more solid theoretical foundation (De Bortoli et al., 2021).
Based on these progresses, SGMs and related models show great potential in various application(Yang et al., 2022, 2023) including high fidelity audio streams generation (Kong et al., 2021; Chen et al., 2021), text-to-image generation (Nichol et al., 2021; Saharia et al., 2022; Rombach et al., 2022), photo-realistic images editing (Meng et al., 2021; Saharia et al., 2021a, b; Kawar et al., 2022), video (Ho et al., 2022) generation, 3D shape generation (Zhou et al., 2021), and scene graphs generation (Yang et al., 2022). Starting from a sample initialized with a Gaussian distribution, an SGM produces a target sample by simulating a sampling process, typically a reverse diffusion, Langevin dynamics, or their combination (Alg. 1). Compared to state-of-the-art GANs (Brock et al., 2019; Karras et al., 2019, 2018), a significant drawback of existing SGMs is drastically slower generation due to the need to take many iterations for a sequential sampling process (Song et al., 2021b; Nichol et al., 2021; Xiao et al., 2022).
@@
Conceptually, the Langevin dynamics defined in Eq. (1) can be considered as a special case of Metropolis adjusted Langevin algorithm (MALA) (Roberts and Stramer, 2002; Welling and Teh, 2011; Girolami and Calderhead, 2011). When the coordinates of the target dataset (e.g., the pixel locations of a natural image) have widely varying scales or are strongly correlated, the sampling process becomes ill-conditioned. As a result, the convergence of the sampling process is not guaranteed when the step sizes are large (Girolami and Calderhead, 2011). This explains the failure of sampling under direct acceleration (Fig. 1). Noting that the reverse diffusion has a similar update rule to the Langevin dynamics (Eq. (1) vs. Eq. (3)), we hypothesize that it suffers from the same ill-conditioned issue.
—————–$T=2000$———————$T=200$———————$T=133$———————$T=100$———————$T=66$
——————    Ours               Baseline Song et al. (2021b)
Figure 1: Facial images at a resolution of $1024\times 1024$ generated by NCSN++ (Song et al., 2021b) under a variety of sampling iterations (top) without and (bottom) with our PDS. NCSN++ decades quickly with increasingly reduced sampling iterations, which can be well solved with PDS. For generating a batch of 8 images, PDS reduces the time cost of generating $8$ human facial samples from $1920$ sec (more than half an hours) to $68$ sec (about one minute) on one NVIDIA RTX 3090 GPU. Dataset: FFHQ (Karras et al., 2019). More samples in Appendix.
@@
In light of this insight, we propose an efficient preconditioned diffusion sampling (PDS) method for accelerating existing pretrained SGMs, without the need for model retraining. The key idea is to make the scales of the target distribution more similar along all the coordinates (Roberts and Stramer, 2002; Li et al., 2016) using a matrix preconditioning operation, hence solving the ill-conditioned problem of both the Langevin dynamics Eq. (1) and the reverse diffusion Eq. (3), as demonstrated in Fig. 2. Formally, we enrich the Langevin dynamics Eq. (1) by reformulating the sampling process as
@@
Theoretically, we prove that PDS can preserve both the steady-state distribution of the original Langevin dynamics and the final-state distribution of the original reverse diffusion, with the assistance of Fokker-Planck equation (Gardiner et al., 1985) and related techniques. We construct the preconditioning operator $M$ using Fast Fourier Transform (FFT) (Brigham, 1988), making its computational cost marginal.
In this work, we make the following contributions: (1) For sampling acceleration, we introduce a novel preconditioned diffusion sampling (PDS) process. PDS preconditions the existing sampling process for alleviating ill-conditioned issues, with a theoretical guarantee of keeping the original target distributions in convergence. (2) With PDS, various pretrained SGMs can be accelerated significantly for image synthesis at various spatial resolutions, without model retraining. In particular, PDS delivers 28×28$\times$28 × reduction in wall-clock time for high-resolution $(1024\times 1024)$ image synthesis. Compared with latest alternatives (e.g., CLD-SGM (Dockhorn et al., 2022), DDIM (Song et al., 2020), and Analytic-DDIM (Bao et al., 2022b)), PDS achieves superior performance on CIFAR-10 (FID score 1.99).
A preliminary version of this work was presented in ECCV 2022 (Ma et al., 2022). In this paper, we have further extended our conference version as follows: (i) We reveal that the sampling process of the original SGMs is insensitive to the structural data (Sec. 3.4); (ii) We provide more theoretical explanation including its effect of preserving the final-state distribution of the reverse diffusion during sampling (Sec. 4.3); (iii) We develop indicators that further illustrate the mechanism of the PDS in terms of alleviating the ill-conditioning issue. (iv) We conduct more quantitative experiments, including an extra dataset (CelebA, COCO and ImageNet), more iteration budgets evaluation, and comparing with more acceleration methods in both SGMs and DDPMs; (v) We investigate the parameter space of PDS more comprehensively, resulting in further performance improvement; (vi) We develop a theory easing the process of estimating proper parameters (Sec. 6.3) with experimental validation (Sec. 8.2). (vii) We implement sensitivity analysis on the parameter of PDS and reveal its robustness.
@@ 2 Related work
Sohl-Dickstein et al. (2015) first proposed to sequentially corrupt a target data distribution with Gaussian noises and train a model to learn the backward process to recover the target data distribution, inspired by non-equilibrium statistical physics. Exploring this procedure later on, Song and Ermon (2019) proposed the first SGM, the noise conditional score network (NCSN). They further proposed NCSNv2 for higher resolution image generation (e.g., $256\times 256$) by scaling noises and improving stability with moving average Song and Ermon (2020). Summarizing all the previous SGMs into a unified framework based on the stochastic differential equation (SDE), Song et al. (2021b) proposed NCSN++ to generate high-resolution images via numerical SDE solvers for the first time. De Bortoli et al. (2021) provided the first quantitative convergence results for SGMs. As another class of relevant generative models, denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020; Dhariwal and Nichol, 2021; Ho et al., 2021; Nichol et al., 2021) also demonstrate excellent performance on image synthesis. They are mostly trained by maximizing an evidence lower bound (ELBO).
@@
There are some recent works on improving the sampling quality of SGMs and DDPMs by reshaping the diffusion process. Song et al. (2020) proposed DDIMs by improving the sampling of DDPMs via a non-Markovian sampling process. Liu et al. (2021) further improved DDIMs by implementing the sampling process near a manifold to guarantee that the model can fit the ground-truth score function well. Using the Hamiltonian Monte Carlo methods (Neal et al., 2011), Dockhorn et al. (2022) extended the sample space of SGMs with a velocity vector, and proposed critically-damped Langevin diffusion (CLD) based SGMs. Jolicoeur-Martineau et al. (2021) used a numerical SDE solver with adaptive step sizes to accelerate SGMs.
@@
As our PDS, some prior methods utilize non-isotropic noises to accelerate the sampling process. Jing et al. (2022) restricted the sampling process in a linear subspace where the data reside to accelerate the convergence of the sampling process. This might help with the ill-condition issue and convergence speed, as the space of sampling decreases. However, our PDS directly adjusts the scales of different coordinates of the sampling space to alleviate ill-condition issues, without the need for subspace restriction. Nichol and Dhariwal (2021) learned an additional variance vector during training a DDPM. Instead, we exploit a preconditioning matrix simply estimated from the training data, without training extra parameters. With diagonal covariance matrices, Bao et al. (2022b, a) given the analytical formulation of a sequential of diagonal covariance matrices maximizing the ELBO between the corrupting process and the sampling process for better sample performance. In comparison, our method eliminates the need for solving additional complex optimization problems.
Some works focus on improving the sampling efficiency of SGMs or DDPMs by incorporating extra learnable modules. Latent Diffusion Models (LDMs) use a nonlinear transformation induced by a variational autoencoder to transfer the sampling process of SGMs into a latent space (Vahdat et al., 2021; Rombach et al., 2022). The lower-dimensional, decorrelated nature of this latent space helps mitigate ill-conditioned issues similar to PDS. However, unlike LDM, our PDS employs a train-free linear transform, offering a more cost-effective solution. A number of recent works (Salimans and Ho, 2022; Meng et al., 2023; Song et al., 2023) utilized the concept of knowledge distillation to expedite the sampling process of DDPMs. Their common idea is a progressive distillation approach, involving the iterative refinement of a pretrained model. Kingma et al. (2021) proposed variations diffusion models to learn the noise schedule during training with another neural network for higher log-likelihood of the sampling results. In comparison, our PDS directly leverages information from the data, avoiding the need for extra model training to accelerate sampling
Efforts to enhance the sampling quality of SGMs are also being made by modifying their training loss. Song et al. (2021a) proposed reweighting the loss function with the diffusion coefficient of the SGMs. This enhancement aims to boost the log-likelihood of the model’s generations, effectively improving the performance of VP (variance preserving) and subVP SGMs. However, this method cannot yield improvements for VE (variance exploding) SGMs. Further development in this direction is (Lu et al., 2022), where they proposed training the model to minimize higher-order score matching errors. This technique notably improves the performance of ordinary differential equation (ODE) type SGMs. In comparison, our PDS demonstrates superior improvement on the stochastic differential equation (SDE) type VE (variance exploding) SGMs.
@@
The above methods are limited in the following aspects: (1) They tend to introduce much extra computation. For example, CLD-SGMs double the dimension of data by learning the velocity of diffusion. Jolicoeur-Martineau et al. (2021) added a high-order numerical solver that increases the number of calling the SGM, resulting in much more cost. With our PDS, the only extra calculation relates the preconditioning matrix that can be efficiently implemented by Fast Fourier Transform. As a result, PDS yields marginal extra computation to the vanilla SGMs. (2) They often improve a single specific SGM/DDPM while our PDS is able to accelerate three different SGMs (NCSN (Song and Ermon, 2019), NCSNv2 (Song and Ermon, 2020), and NCSN++ (Song et al., 2021b)). (3) Unlike our work, none of the previous methods has demonstrated the scalability to more challenging high-resolution image generation tasks (e.g., FFHQ facial images with $1024\times 1024$ resolution).
@@ 3.1 Score matching
In the absence of $Z(\theta)$, training with this objective is more tractable. Hyvärinen and Dayan (2005) further proves that under some mild conditions, if $p_{\mathbf{x}^{\ast}}(\cdot)=p_{\mathbf{x}^{\ast}}(\cdot;\theta^{\ast})$, the ground-truth parameters $\theta^{\ast}$ can be achieved by minimizing Eq. (6) to zero.
However, Eq. (6) can not be evaluated in more general cases where an expression of the score function about $\theta$ is absent. To address this, Vincent (2011) proposes an equivalent, simpler objective inspired by denoising autoencoders
@@
where $p_{\sigma}(\mathbf{x}_{\sigma}|\mathbf{x})=\mathcal{N}(\mathbf{x}_{\sigma};\mathbf{x},\sigma^{2}I)$. Formally, given a noised data sample $\mathbf{x}_{\sigma}$, the model $\mathbf{s}_{\theta}(\cdot)$ is trained to extract the Gaussian noise vector added into the sample. This denoises the sample corrupted by the noise, i.e., subtracting the noise vector gives the original sample. It is proved that minimizing the objective Eq. (7) allows the model to fit the ground-truth score function $\mathbf{s}_{\mathbf{x}^{\ast}}$ of the data distribution (Vincent, 2011).
3.2 Langevin dynamics
SGMs generate samples from a target distribution by iteratively evaluating the target score function. The first SGM, noise conditional score network (NCSN) (Song and Ermon, 2019) $\mathbf{s}_{\theta}$, is trained on a series of score matching tasks via
$\sum_{t=1}^{T}\sigma^{2}_{t}\mathbb{E}_{\mathbf{x}\sim p_{\mathbf{x}}}\left[\left\|\mathbf{s}_{\theta}(\mathbf{x}_{\sigma_{t}},\sigma_{t})-\bigtriangledown_{\mathbf{x}_{\sigma_{t}}}\log p_{\sigma_{t}}(\mathbf{x}_{\sigma_{t}}|\mathbf{x})\right\|^{2}\right],$
with the standard deviations $\{\sigma_{t}\}_{t=1}^{T}$ satisfies $0<\sigma_{1}<\sigma_{2}<\ldots<\sigma_{T}$. During inference, an NCSN generates samples from $p_{\mathbf{x}^{\ast}}$ through the discretized Langevin dynamics